A metal worker has a metal alloy that is 30% copper and another alloy that is 80% copper. How many kilograms of each alloy should the metalworker combine to creat 60 kilograms of a 51% copper alloy?
Answer by lwsshak3(11628)

A metal worker has a metal alloy that is 30% copper and another alloy that is 80% copper. How many kilograms of each alloy should the metalworker combine to creat 60 kilograms of a 51% copper alloy?
let x=amt of 30% copper alloy to combine
60-x=amt of 80% copper alloy to combine
..
30%x+80%(60-x)=51%*60
.30x+40-.80x=30.6
.50x=9.4
x=50.6
How many kilograms of each alloy should the metalworker combine?
30% copper alloy: 9.4 kg
80% copper alloy: 50.6 kg